If you can drive 40 miles on side roads in the same amount of time as it takes to drive 30 miles on the highway. If the rate of speed is 7mph faster on the back roads then the highway, what is the rate of speed on the side roads?
Answer by Nate(3500)

distance / rate = time
Highway:
distance = 30
rate = x
time = 30 / x
Side Roads:
distance = 40
rate = x + 7
time = 40 / (x + 7)
Time were the same:
30 / x = 40 / (x + 7)
30(x + 7) = 40x
30x + 210 = 40x
210 = 10x
21 = x
What is the rate of speed on the side roads?
28 mph